# PAC in DESI. I. Galaxy Stellar Mass Function into the $10^{6}{\rm M}_{\odot}$ Frontier

### v2

@@ 2.2 DECaLS
Image processing was performed using Tractor (Lang et al., 2016) to extract sources. These sources are modelled using parametric profiles convolved with specific point spread functions (PSFs), including a delta function for point sources, an exponential profile, a de Vaucouleurs profile, and a Sérsic profile. The best-fit model magnitudes are used as the default photometry, with correction for Galactic extinction (Schlegel et al., 1998).
@@
Figure 2: $V_{\rm{max}}$-weighted rest-frame $g-r$ colour vs. stellar mass distribution for BGS with $z<0.2$. The distribution is further normalized within each stellar mass bin for clearer visualization. The line $g-r=0.04\log_{10}(M_{*}/{\rm M}_{\sun})+0.2$ is used to separate the blue and red populations.
Figure 3: Mean errors and 1 $\sigma$ scatter in interpolated stellar mass $\Delta\log_{10}M_{*}$ at 241 redshifts interpolated from the values at 44 reference redshifts within the range of $10.001<z<$ 1. For these reference redshifts, 31 of them are chosen with an equal interval of $\Delta\log_{10}(1+z)=0.01$, while the other 13 indicated by dotted lines are included to account for the rapid changes in stellar mass with a given redshift due to specific structures in the galaxy spectrum.
Figure 4: The observed $g-r$ vs. $r-z$ colour diagrams for DESI BGS for two redshift bins. The red solid lines represent the colour cuts used to exclude background galaxies in DECaLS for PAC measurements at each redshift bin.
@@ 3.1 Overview of the previous PAC method
Suppose we aim to investigate two distinct populations of galaxies: one, denoted as ${\rm{pop}}_{1}$, is drawn from a spectroscopic catalogue, while the other, referred to as ${\rm{pop}}_{2}$, is sourced from a photometric catalogue. These populations are characterized within a relatively narrow redshift range. In our prior work Paper I, we introduced a method called PAC, which enables the accurate measurement of the excess surface density, $\bar{n}_{2}w_{{\rm{p}}}(r_{\rm{p}})$, of objects with selected physical properties from ${\rm{pop}}_{2}$ around objects from ${\rm{pop}}_{1}$. It is based on
@@
For more details, see Paper I.
With this method, $\bar{n}_{2}w_{{\rm{p}}}$ is calculated in each redshift bin and then combined, potentially introducing two sources of error. First, the properties of ${\rm{pop}}_{2}$ are all computed using the mean redshift of the bin, while they are correlated with ${\rm{pop}}_{1}$ objects spanning the full width of the redshift bin. This can lead to inaccuracies in the properties due to differences in luminosity distances and the rest-frame wavelength covered by the photometric bands. Moreover, the measured $\omega_{12,\rm{w}}(\theta)$ is multiplied by the same $\bar{S}_{2}/r_{1}^{2}$ factor, even though it should vary for each source in ${\rm{pop}}_{1}$. Although, as demonstrated in Paper I, these approximations are found to be acceptable at high redshift, they become less accurate at lower redshifts ($z<0.1$) due to the much more significant fractional variation in $r_{1}$. Hence, we now present an improved PAC method that operates without the need for redshift bins.
@@ 3.2 Properties of photometric objects across the entire redshift range
We perform SED fitting to the 1 per cent sub-sample at 241 redshifts equally spaced in $\Delta\log_{10}(1+z)$ within the redshift range of $10.001<z<$ 1. As demonstrated in Figure 3, our results indicate that employing 31 reference redshifts with spacing $\Delta\log_{10}(1+z)=0.01$, in combination with an extra 13 redshifts linked to specific features in galaxy spectra, enables accurate interpolation of the stellar masses for all sources throughout the entire redshift range. The need for the 13 extra redshifts can be illustrated by considering the case of $z\sim 0.13$ where the D4000 feature begins to enter the $g$ band. Here, many galaxies that at a lower assumed redshift would be classified as red are suddenly classified as blue. This results in the inferred stellar mass varying rapidly with redshift, necessitating the inclusion of additional reference redshifts to maintain the same level of accuracy at these specific redshifts. Using these 44 reference redshifts, we perform the interpolation under the assumption that $\log_{10}(M_{*})$ varies linearly with $\log_{10}(D_{\rm{L}})$, where $D_{\rm{L}}$ denotes the luminosity distance. This approach achieves an accuracy of $\Delta\log_{10}M_{*}<0.02$, validated by comparing interpolated $M_{*}$ values with those directly calculated on a finer grid, as shown in Figure 3.
@@
Subsequently, we perform SED fitting for the entire DECaLS sample at the 44 reference redshifts. While it is possible to obtain the properties at the redshift of each ${\rm{pop}}_{1}$ source, such an approach is unrealistic and unnecessary. Instead, we opt for interpolating and saving the stellar masses at 634 redshifts within the range of $10.001<z<$ 1. The mean change in stellar mass between adjacent redshifts in this interpolated dataset is smaller than $\Delta\log_{10}M_{*}=0.01$, which provides a level of accuracy that is sufficient for our analysis. For example, if 2 adjacent reference redshifts among the 44 identified above show a mean $\Delta\log_{10}M_{*}=0.1$, we insert 9 additional redshifts in between, ensuring a consistent spacing of $\Delta\log_{10}M_{*}$ values through interpolation. For each ${\rm{pop}}_{1}$ source, we utilize the physical properties of ${\rm{pop}}_{2}$ at the nearest available redshift.
Although the above approach is illustrated using stellar mass as an example, it can be applied to any other physical properties under investigation. We have tested the precision of the interpolation for rest-frame $g$- and $r$-band luminosities and found similar accuracy. Consequently, we expect that the precision for the rest-frame $g-r$ colour is also sufficient for our analysis.
@@ 3.3 A new estimator for the PAC method
In the previous PAC method, for each redshift bin we first compute $\omega_{12,\rm{w}}$ using the Landy–Szalay estimator (Landy & Szalay, 1993) with two corresponding random samples
@@
where $DD$, $DR$ and $RR$ are the data-data, data-random and random-random pair counts. The angle $\theta$ varies for different pairs of $D_{1}$ and $R_{1}$ and is defined as $\theta=\arcsin{(r_{\rm{p}}/r_{1})}$. Subsequently, $\bar{n}_{2}w_{{\rm{p}}}(r_{\rm{p}})$ is obtained using Equation 2 by multiplying it with $\bar{S}_{2}/r_{1}^{2}$, which is calculated using the mean redshift of each bin.
@@ 3.4 Excluding foreground and background galaxies with colour cuts
As an example, in this paper, we focus on the redshift range of interest at $z<0.2$. In Figure 4, we plot the DESI Y1 BGS galaxies in two narrower redshift bins $z<0.1$ and $0.20.1<z<$ 0.2 on the $g-r$ vs. $r-z$ colour-colour diagram. According to their colour distribution, we derived colour cuts for these two redshift ranges. For $z<0.1$,
@@
The $g-r$ color cut range we used is sufficiently wide compared to the color distribution of BGS in Figure 2, ensuring that all low-redshift sources are retained. These colour cuts result in a reduction of more than half the galaxies in DECaLS. Additionally, it is important to note that $\bar{S}_{2}$ should also be calculated using the sample after applying the colour cut, rather than the entire sample. This approach can be extended to higher redshifts with an appropriate reference sample, which will be explored in future studies.
@@ 3.5 Determining stellar mass limits
[…]
@@
We repeat the process at various redshifts for $z<0.2$ for the entire, blue, and red populations. We present the results in the top panel of Figure 6 and also plot the results for the entire population from Paper III based on GAMA data for comparison. The results are discrete because we measure the angular surface density in stellar mass bins with an interval of $\Delta\log_{10}(M_{*}/{\rm M}_{\sun})=0.2$. The stellar mass limits are nearly identical for the red and entire populations but are lower for blue galaxies. This is because red galaxies have higher mass-to-light ratios than blue ones, making them fainter at fixed stellar masses. As a result, the stellar mass limits for the entire population are primarily determined by the red galaxies. Compared to Paper III, we find that the stellar mass limits derived from GAMA are overly optimistic. This may suggest that GAMA does not fully sample the entire colour space of low-mass galaxies, potentially missing red ones. In the bottom panel, we further display the effective volume calculated using the maximum complete redshift, as well as with a redshift cutoff of $z>0.01$, for various stellar masses. Our results indicate the potential to investigate galaxies with stellar masses of $10^{6}{\rm M}_{\sun}$ within a sufficiently large volume using the PAC method with BGS and DECaLS.
@@
When utilizing low redshift data, it becomes necessary to re-evaluate the approximation in Equation 2. Wang et al. (2011) has demonstrated an accuracy of 3 per cent even at $r_{\rm{p}}=30h^{-1}\rm{Mpc}$ for various $r$-band absolute magnitude bins at $z=0.07$. We intend to expand the analysis to lower redshifts, encompass a broader range of galaxy populations, and incorporate additional sources of error that were not considered in Wang et al. (2011).
The first error arises from the different definitions of $w_{{\rm{p}}}(r_{\rm{p}})$ and $\omega_{12}(\theta)$ that was comprehensively addressed in Wang et al. (2011). Assuming we have a ${\rm{pop}}_{1}$ sample at a specific redshift, by definition, we have
@@
However, there is another source of error that has not been considered in Wang et al. (2011). Suppose we are interested in obtaining measurements for ${\rm{pop}}_{2}$ within a specific stellar mass bin $[M_{\rm{min}},M_{\rm{max}}]$ and the stellar masses of ${\rm{pop}}_{2}$ are calculated at the redshift of ${\rm{pop}}_{1}$ as described in Section 3.2. When we compute the angular correlation, ${\rm{pop}}_{1}$ is correlated with some galaxies with $M_{*}>M_{\rm{max}}$ at $r_{2}>r_{1}$ and $M_{*}<M_{\rm{min}}$ at $r_{2}<r_{1}$ because the masses of correlated galaxies are either underestimated or overestimated due to the inaccurate redshift assumed. This effect can impact the measurements because both $\bar{n}_{2}$ and $\xi(r_{12})$ depend on stellar mass, which is also negligible when $r_{1}$ is significantly larger than the correlation length. To investigate this effect, we re-write the $\omega_{12}(\theta)$ we measured as
@@
To estimate the accuracy of the approximation $\Delta\eta/\eta_{\rm{true}}$ for various redshifts and stellar mass bins, we employ the precise stellar-halo mass relation (SHMR) from Paper IV and the CosmicGrowth $N$-body simulation (Jing, 2019) with a subhalo catalogue constructed using HBT+ (Han et al., 2012, 2016). The CosmicGrowth simulation has a mass resolution of $5\times 10^{8}h^{-1}{\rm M}_{\sun}$ and a box size of $600\,h^{-1}\,\text{Mpc}$, which is appropriate for studying the distribution of galaxies with $M_{*}>10^{8}{\rm M}_{\sun}$ according to Paper IV. We derive $\xi(r,M_{*})$ and $\Phi(M_{*})$ using the SHMR and simulation data, and then estimate $\Delta\eta/\eta_{\rm{true}}$ across different redshifts and for varying stellar masses of ${\rm{pop}}_{1}$ and ${\rm{pop}}_{2}$. Figure 7 presents the results at 2 characteristic scales, $r_{\rm{p}}=1.25h^{-1}\rm{Mpc}$ and $12.5h^{-1}\rm{Mpc}$, across 5 redshifts $z=0.01,0.025,0.075$, and $0.1$, and within 18 stellar mass bins for ${\rm{pop}}_{1}$ and ${\rm{pop}}_{2}$, spanning the range from $10^{8.0}{\rm M}_{\sun}$ to $10^{11.6}{\rm M}_{\sun}$. We find that $\Delta\eta/\eta_{\rm{true}}$ is smaller for smaller scales, lower ${\rm{pop}}_{2}$ mass, and higher redshift, with its dependence on the mass of ${\rm{pop}}_{1}$ being nearly negligible. There are indications that $\Delta\eta/\eta_{\rm{true}}$ begins to increase when $M_{\rm{pop}_{2}}<10^{9.0}{\rm M}_{\sun}$. This behaviour is attributed to the fact that within these mass bins, ${\rm{pop}}_{1}$ starts to correlate with some galaxies having $M_{\rm{pop}_{2}}<10^{8.0}{\rm M}_{\sun}$, whose host halos or subhalos are unresolved by the simulation. In general, we can maintain an accuracy of better than 3 per cent for almost all the stellar mass bins at $z=0.05$ within a scale of up to $12.5h^{-1}\rm{Mpc}$. It is worth noting that the measurements become less accurate in the most massive ${\rm{pop}}_{2}$ bins at lower redshifts. However, these redshifts are primarily intended for the study of low mass galaxies and have fewer massive galaxies. The PAC method remains highly accurate for $M_{\rm{pop}_{2}}<10^{9.0}{\rm M}_{\sun}$ even at $z=0.01$, and we anticipate that this level of accuracy extends to lower mass galaxies. Although the test above was conducted for the entire sample, we expect that the conclusion also holds for the red and blue sub-samples. This demonstrates the potential of the PAC method for exploring the unknown region of faint galaxies.
@@
Figure 10: Comparison of $w_{{\rm{p}}}$ at $8h^{-1}\rm{Mpc}$ for red ${\rm{pop}_{2}^{\rm{s}}}$ across various stellar mass bins. Each panel corresponds to one of the 11 ${\rm{pop}_{1}}$ stellar mass bins. The orange solid lines with shaded regions represent the results from the $w_{{\rm{p}}}$ of all ${\rm{pop}_{2}^{\rm{s}}}$ with $M_{*}<10^{10.8}{\rm M}_{\sun}$. For comparison, the projected matter correlations $w_{{\rm{p}}}^{\rm{mm}}$ at redshifts 0 (dashed lines) and 0.1 (dotted lines) calculated using CAMB (Lewis et al., 2000) are also shown.

### v3

@@ 2.2 DECaLS
Image processing was performed using Tractor (2016ascl.soft04008L) to extract sources. These sources are modelled using parametric profiles convolved with specific point spread functions (PSFs), including a delta function for point sources, an exponential profile, a de Vaucouleurs profile, and a Sérsic profile. The best-fit model magnitudes are used as the default photometry, with correction for Galactic extinction (1998ApJ...500..525S).
@@
Figure 2: $V_{\rm{max}}$-weighted rest-frame grg-r colour vs. stellar mass distribution for BGS with $z<0.2$. The distribution is further normalized within each stellar mass bin for clearer visualization. The line $g-r=0.04\log_{10}(M_{*}/{\rm M}_{\sun})+0.2$ is used to separate the blue and red populations.
Figure 3: Mean errors and 1 $\sigma$ scatter in interpolated stellar mass $\Delta\log_{10}M_{*}$ at 241 redshifts interpolated from the values at 44 reference redshifts within the range of $0.001<z<1$. For these reference redshifts, 31 of them are chosen with an equal interval of $\Delta\log_{10}(1+z)=0.01$, while the other 13 indicated by dotted lines are included to account for the rapid changes in stellar mass with a given redshift due to specific structures in the galaxy spectrum.
Figure 4: The observed grg-r vs. rzr-z colour diagrams for DESI BGS for two redshift bins. The red solid lines represent the colour cuts used to exclude background galaxies in DECaLS for PAC measurements at each redshift bin.
@@ 3.1 Overview of the previous PAC method
Suppose we aim to investigate two distinct populations of galaxies: one, denoted as ${\rm{pop}}_{1}$, is drawn from a spectroscopic catalogue, while the other, referred to as ${\rm{pop}}_{2}$, is sourced from a photometric catalogue. These populations are characterized within a relatively narrow redshift range. In our prior work 2022ApJ...925...31X, we introduced a method called PAC, which enables the accurate measurement of the excess surface density, $\bar{n}_{2}w_{{\rm{p}}}(r_{\rm{p}})$, of objects with selected physical properties from ${\rm{pop}}_{2}$ around objects from ${\rm{pop}}_{1}$. It is based on
@@
For more details, see 2022ApJ...925...31X.
With this method, $\bar{n}_{2}w_{{\rm{p}}}$ is calculated in each redshift bin and then combined, potentially introducing two sources of error. First, the properties of ${\rm{pop}}_{2}$ are all computed using the mean redshift of the bin, while they are correlated with ${\rm{pop}}_{1}$ objects spanning the full width of the redshift bin. This can lead to inaccuracies in the properties due to differences in luminosity distances and the rest-frame wavelength covered by the photometric bands. Moreover, the measured $\omega_{12,\rm{w}}(\theta)$ is multiplied by the same $\bar{S}_{2}/r_{1}^{2}$ factor, even though it should vary for each source in ${\rm{pop}}_{1}$. Although, as demonstrated in 2022ApJ...925...31X, these approximations are found to be acceptable at high redshift, they become less accurate at lower redshifts ($z<0.1$) due to the much more significant fractional variation in $r_{1}$. Hence, we now present an improved PAC method that operates without the need for redshift bins.
@@ 3.2 Properties of photometric objects across the entire redshift range
We perform SED fitting to the 1 per cent sub-sample at 241 redshifts equally spaced in $\Delta\log_{10}(1+z)$ within the redshift range of $0.001<z<1$. As demonstrated in Figure 3, our results indicate that employing 31 reference redshifts with spacing $\Delta\log_{10}(1+z)=0.01$, in combination with an extra 13 redshifts linked to specific features in galaxy spectra, enables accurate interpolation of the stellar masses for all sources throughout the entire redshift range. The need for the 13 extra redshifts can be illustrated by considering the case of $z\sim 0.13$ where the D4000 feature begins to enter the $g$ band. Here, many galaxies that at a lower assumed redshift would be classified as red are suddenly classified as blue. This results in the inferred stellar mass varying rapidly with redshift, necessitating the inclusion of additional reference redshifts to maintain the same level of accuracy at these specific redshifts. Using these 44 reference redshifts, we perform the interpolation under the assumption that $\log_{10}(M_{*})$ varies linearly with $\log_{10}(D_{\rm{L}})$, where $D_{\rm{L}}$ denotes the luminosity distance. This approach achieves an accuracy of $\Delta\log_{10}M_{*}<0.02$, validated by comparing interpolated $M_{*}$ values with those directly calculated on a finer grid, as shown in Figure 3.
@@
Subsequently, we perform SED fitting for the entire DECaLS sample at the 44 reference redshifts. While it is possible to obtain the properties at the redshift of each ${\rm{pop}}_{1}$ source, such an approach is unrealistic and unnecessary. Instead, we opt for interpolating and saving the stellar masses at 634 redshifts within the range of $0.001<z<1$. The mean change in stellar mass between adjacent redshifts in this interpolated dataset is smaller than $\Delta\log_{10}M_{*}=0.01$, which provides a level of accuracy that is sufficient for our analysis. For example, if 2 adjacent reference redshifts among the 44 identified above show a mean $\Delta\log_{10}M_{*}=0.1$, we insert 9 additional redshifts in between, ensuring a consistent spacing of $\Delta\log_{10}M_{*}$ values through interpolation. For each ${\rm{pop}}_{1}$ source, we utilize the physical properties of ${\rm{pop}}_{2}$ at the nearest available redshift.
Although the above approach is illustrated using stellar mass as an example, it can be applied to any other physical properties under investigation. We have tested the precision of the interpolation for rest-frame $g$- and $r$-band luminosities and found similar accuracy. Consequently, we expect that the precision for the rest-frame grg-r colour is also sufficient for our analysis.
@@ 3.3 A new estimator for the PAC method
In the previous PAC method, for each redshift bin we first compute $\omega_{12,\rm{w}}$ using the Landy–Szalay estimator (1993ApJ...412...64L) with two corresponding random samples
@@
where $DD$, DRDR and $RR$ are the data-data, data-random and random-random pair counts. The angle $\theta$ varies for different pairs of $D_{1}$ and $R_{1}$ and is defined as $\theta=\arcsin{(r_{\rm{p}}/r_{1})}$. Subsequently, $\bar{n}_{2}w_{{\rm{p}}}(r_{\rm{p}})$ is obtained using Equation 2 by multiplying it with $\bar{S}_{2}/r_{1}^{2}$, which is calculated using the mean redshift of each bin.
@@ 3.4 Excluding foreground and background galaxies with colour cuts
As an example, in this paper, we focus on the redshift range of interest at $z<0.2$. In Figure 4, we plot the DESI Y1 BGS galaxies in two narrower redshift bins $z<0.1$ and $0.1<z<0.2$ on the grg-r vs. rzr-z colour-colour diagram. According to their colour distribution, we derived colour cuts for these two redshift ranges. For $z<0.1$,
@@
The grg-r color cut range we used is sufficiently wide compared to the color distribution of BGS in Figure 2, ensuring that all low-redshift sources are retained. These colour cuts result in a reduction of more than half the galaxies in DECaLS. Additionally, it is important to note that $\bar{S}_{2}$ should also be calculated using the sample after applying the colour cut, rather than the entire sample. This approach can be extended to higher redshifts with an appropriate reference sample, which will be explored in future studies.
@@ 3.5 Determining stellar mass limits
While it is evident that DECaLS can extend to fainter magnitude and lower mass limits, determining these values is essential for fully exploiting the dataset. However, this task is not straightforward as we are reaching unexplored territory, with limited knowledge about these faint galaxies. In our previous studies (2022ApJ...939..104X; 2023ApJ...944..200X), we first determined the $r$ or $z$ band galaxy depths of DECaLS. Then, we created apparent magnitude vs. stellar mass diagrams using GAMA DR4 data (2022MNRAS.513..439D) with $V_{\rm{max}}$ correction for galaxies at $z<0.2$ and using DES deep field data (2022MNRAS.509.3547H) with photoz for galaxies at $0.2<z<0.7$. Subsequently, we determined the stellar mass for which DECaLS achieves 95 per cent completeness at each redshift with the galaxy depths we obtained. These results are sensitive to the determination of galaxy depths, as well as to corrections for incompleteness, cosmic variance in GAMA, and photo-z errors in DES. Additionally, both GAMA and DESI struggle to measure enough galaxies with $M_{*}<10^{8.0}{\rm M}_{\sun}$, making this method impractical for the lower mass end.
@@
We repeat the process at various redshifts for $z<0.2$ for the entire, blue, and red populations. We present the results in the top panel of Figure 6 and also plot the results for the entire population from 2022ApJ...939..104X based on GAMA data for comparison. The results are discrete because we measure the angular surface density in stellar mass bins with an interval of $\Delta\log_{10}(M_{*}/{\rm M}_{\sun})=0.2$. The stellar mass limits are nearly identical for the red and entire populations but are lower for blue galaxies. This is because red galaxies have higher mass-to-light ratios than blue ones, making them fainter at fixed stellar masses. As a result, the stellar mass limits for the entire population are primarily determined by the red galaxies. Compared to 2022ApJ...939..104X, we find that the stellar mass limits derived from GAMA are overly optimistic. This may suggest that GAMA does not fully sample the entire colour space of low-mass galaxies, potentially missing red ones. In the bottom panel, we further display the effective volume calculated using the maximum complete redshift, as well as with a redshift cutoff of $z>0.01$, for various stellar masses. Our results indicate the potential to investigate galaxies with stellar masses of $10^{6}{\rm M}_{\sun}$ within a sufficiently large volume using the PAC method with BGS and DECaLS.
@@
When utilizing low redshift data, it becomes necessary to re-evaluate the approximation in Equation 2. 2011ApJ...734...88W has demonstrated an accuracy of 3 per cent even at $r_{\rm{p}}=30h^{-1}\rm{Mpc}$ for various $r$-band absolute magnitude bins at $z=0.07$. We intend to expand the analysis to lower redshifts, encompass a broader range of galaxy populations, and incorporate additional sources of error that were not considered in 2011ApJ...734...88W.
The first error arises from the different definitions of $w_{{\rm{p}}}(r_{\rm{p}})$ and $\omega_{12}(\theta)$ that was comprehensively addressed in 2011ApJ...734...88W. Assuming we have a ${\rm{pop}}_{1}$ sample at a specific redshift, by definition, we have
@@
However, there is another source of error that has not been considered in 2011ApJ...734...88W. Suppose we are interested in obtaining measurements for ${\rm{pop}}_{2}$ within a specific stellar mass bin $M_{\rm{min}},M_{\rm{max}}]$ and the stellar masses of ${\rm{pop}}_{2}$ are calculated at the redshift of ${\rm{pop}}_{1}$ as described in Section 3.2. When we compute the angular correlation, ${\rm{pop}}_{1}$ is correlated with some galaxies with $M_{*}>M_{\rm{max}}$ at $r_{2}>r_{1}$ and $M_{*}<M_{\rm{min}}$ at $r_{2}<r_{1}$ because the masses of correlated galaxies are either underestimated or overestimated due to the inaccurate redshift assumed. This effect can impact the measurements because both $\bar{n}_{2}$ and $\xi(r_{12})$ depend on stellar mass, which is also negligible when $r_{1}$ is significantly larger than the correlation length. To investigate this effect, we re-write the $\omega_{12}(\theta)$ we measured as
@@
To estimate the accuracy of the approximation $\Delta\eta/\eta_{\rm{true}}$ for various redshifts and stellar mass bins, we employ the precise stellar-halo mass relation (SHMR) from 2023ApJ...944..200X and the CosmicGrowth $N$-body simulation (2019SCPMA..6219511J) with a subhalo catalogue constructed using HBT+ (2012MNRAS.427.2437H; 2016MNRAS.457.1208H). The CosmicGrowth simulation has a mass resolution of $5\times 10^{8}h^{-1}{\rm M}_{\sun}$ and a box size of $600\,h^{-1}\,\text{Mpc}$, which is appropriate for studying the distribution of galaxies with $M_{*}>10^{8}{\rm M}_{\sun}$ according to 2023ApJ...944..200X. We derive $\xi(r,M_{*})$ and $\Phi(M_{*})$ using the SHMR and simulation data, and then estimate $\Delta\eta/\eta_{\rm{true}}$ across different redshifts and for varying stellar masses of ${\rm{pop}}_{1}$ and ${\rm{pop}}_{2}$. Figure 7 presents the results at 2 characteristic scales, $r_{\rm{p}}=1.25h^{-1}\rm{Mpc}$ and $12.5h^{-1}\rm{Mpc}$, across 5 redshifts $z=0.01,0.025,0.075$, and $0.1$, and within 18 stellar mass bins for ${\rm{pop}}_{1}$ and ${\rm{pop}}_{2}$, spanning the range from $10^{8.0}{\rm M}_{\sun}$ to $10^{11.6}{\rm M}_{\sun}$. We find that $\Delta\eta/\eta_{\rm{true}}$ is smaller for smaller scales, lower ${\rm{pop}}_{2}$ mass, and higher redshift, with its dependence on the mass of ${\rm{pop}}_{1}$ being nearly negligible. There are indications that $\Delta\eta/\eta_{\rm{true}}$ begins to increase when $M_{\rm{pop}_{2}}<10^{9.0}{\rm M}_{\sun}$. This behaviour is attributed to the fact that within these mass bins, ${\rm{pop}}_{1}$ starts to correlate with some galaxies having $M_{\rm{pop}_{2}}<10^{8.0}{\rm M}_{\sun}$, whose host halos or subhalos are unresolved by the simulation. In general, we can maintain an accuracy of better than 3 per cent for almost all the stellar mass bins at $z=0.05$ within a scale of up to $12.5h^{-1}\rm{Mpc}$. It is worth noting that the measurements become less accurate in the most massive ${\rm{pop}}_{2}$ bins at lower redshifts. However, these redshifts are primarily intended for the study of low mass galaxies and have fewer massive galaxies. The PAC method remains highly accurate for $M_{\rm{pop}_{2}}<10^{9.0}{\rm M}_{\sun}$ even at $z=0.01$, and we anticipate that this level of accuracy extends to lower mass galaxies. Although the test above was conducted for the entire sample, we expect that the conclusion also holds for the red and blue sub-samples. This demonstrates the potential of the PAC method for exploring the unknown region of faint galaxies.
@@
Figure 10: Comparison of $w_{{\rm{p}}}$ at $8h^{-1}\rm{Mpc}$ for red $\rm{pop}_{2}^{\rm{s}}$} across various stellar mass bins. Each panel corresponds to one of the 11 $\rm{pop}_{1}$} stellar mass bins. The orange solid lines with shaded regions represent the results from the $w_{{\rm{p}}}$ of all $\rm{pop}_{2}^{\rm{s}}$} with $M_{*}<10^{10.8}{\rm M}_{\sun}$. For comparison, the projected matter correlations $w_{{\rm{p}}}^{\rm{mm}}$ at redshifts 0 (dashed lines) and 0.1 (dotted lines) calculated using CAMB (2000ApJ...538..473L) are also shown.